# footnotetext: * Equal Contribution

### v1

@@ 3.3 Optimization
We prepare the following estimates using off-the-shelf methods in our optimization: 1) masks for the moving objects for each frame $\{\mathbf{M}_{t}\}$, which can be easily obtained using Track-Anything [42, 105] with a few user clicks, 2) monocular depth maps $\{\mathbf{D}_{t}\}$ computed using state-of-the-art relative depth estimation method Depth Anything [106] and 3) long-range 2D tracks $\{\mathbf{U}_{t\rightarrow t^{\prime}}\}$ for foreground pixels from state-of-the-art point tracking method TAPIR [14]. We align the relative depth maps with the metric depth maps by computing a per-frame global scale and shift and use them for our optimization, as we found relative depth maps tend to contain finer details. We treat the lifted 2D tracks unprojected with the aligned depth maps as noisy initial 3D track observations $\{\mathbf{X}_{t}\}$ for the moving objects. For the static part of the scene, we model them using standard static 3D Gaussians and initialize their 3D locations by unprojecting them into the 3D space using the aligned depth maps. The static and dynamic Gaussians are jointly optimized and rasterized together to form an image. We focus on describing the optimization process for dynamic Gaussians below.
@@
Initialization. We first select the canonical frame $t_{0}$ to be the frame in which the most 3D tracks are visible, and initialize the Gaussian means in the canonical frame $\boldsymbol{\mu}_{0}$ as $N$ randomly sampled 3D tracks locations from this set of initial observations. We then perform k-means clustering on the vectorized velocities of the noisy 3D tracks $\{{\mathbf{X}}_{t}\}$, and initialize the motion bases $\{\mathbf{T}_{0\rightarrow t}^{(b)}\}_{b=1}^{B}$ from these $B$ clusters of tracks. Specifically, for the set of trajectories $\{\mathbf{X}_{t}\}_{b}$ belonging to cluster $b$, we initialize the basis transform ${\mathbf{T}}_{0\rightarrow\tau}^{(b)}$ using weighted Procrustes alignment between the point sets $\{\mathbf{X}_{0}\}_{b}$ and $\{\mathbf{X}_{\tau}\}_{b}$ for all $\tau=0,\dots,T$, where the weights are computed using uncertainty and visibility scores from TAPIR predictions. We initialize $\mathbf{w}^{(b)}$ of each Gaussian to decay exponentially with its distance from the center of cluster $b$ in the canonical frame. We then optimize $\boldsymbol{\mu}_{0}$, $\mathbf{w}^{(b)}$, and set of basis functions $\{\mathbf{T}_{0\rightarrow t}^{(b)}\}_{b=1}^{B}$ to fit the observed 3D tracks with an $\ell_{1}$-loss under temporal smoothness constraints.
@@
Training. We supervise the dynamic Gaussians with two sets of losses. The first set of losses comprise our reconstruction loss to match the per-frame pixelwise color, depth, and masks inputs. The second set of losses enforce consistency of correspondences across time. Specifically, during each training step, we render the image $\hat{\mathbf{I}}_{t}$, depth $\hat{\mathbf{D}}_{t}$, and mask $\hat{\mathbf{M}}_{t}$ from their corresponding training cameras $(\mathbf{K}_{t},\mathbf{E}_{t})$ according to Equation 2. We supervise these predictions with a reconstruction loss applied independently per-frame
$L_{\textrm{recon}}=\|\hat{\mathbf{I}}-\mathbf{I}\|_{1}+\lambda_{\text{depth}}\|\hat{\mathbf{D}}-\mathbf{D}\|_{1}+\lambda_{\textrm{mask}}\|\hat{\mathbf{M}}-1\|_{1}.$ (7)
@@
The second set of losses supervises the motion of the Gaussians between frames. Specifically, we additionally render the 2D tracks ${\hat{\mathbf{u}}}_{t\rightarrow t^{\prime}}$ and reprojected depths ${\hat{\mathbf{D}}}_{t\rightarrow t^{\prime}}$, for a pair of randomly sampled query time $t$ and target time $t^{\prime}$. We supervise these rendered correspondences with the lifted long-range 2D track estimates via
$L_{\text{track-2d}}=\|\mathbf{U}_{t\rightarrow t^{\prime}}-\hat{\mathbf{U}}_{t\rightarrow t^{\prime}}\|_{1},~{}\textrm{and}~{}L_{\text{track-depth}}=\|\hat{\mathbf{d}}_{t\rightarrow t^{\prime}}-\hat{\mathbf{D}}(\mathbf{U}_{t\rightarrow t^{\prime}})\|_{1}.$ (8)
@@
Finally, we enforce a distance-preserving loss between randomly sampled dynamic Gaussians and their k-nearest neighbors. Let $\hat{\mathbf{X}}_{t}$ and $\hat{\mathbf{X}}_{t^{\prime}}$ denote the location of a Gaussian at time $t$ and $t^{\prime}$, and $\mathcal{C}_{k}(\hat{\mathbf{X}}_{t})$ denote the set of k-nearest neighbors of $\hat{\mathbf{X}}_{t}$, we define
@@
Implementation Details. For in-the-wild videos, we obtain their camera parameters with the following procedure: we first run UniDepth [66] to get the camera intrinsics and metric depth maps, and then run Droid-SLAM [87] in RGB-D mode with UniDepth’s depth maps to obtain the camera poses. This process is efficient and provides accurate camera parameters. For evaluating our methods on public benchmark, we use the camera annotations that come with the datasets (e.g., from COLMAP [77] or simulation).
We optimize our model using Adam Optimizer [41]. We perform $1000$ iterations of optimization for the initial fitting and $500$ epochs for joint optimization, respectively. The number of $\mathbb{SE}(3)$ bases $B$ is set to 20 for all of our experiments. We initialize $40$k dynamic Gaussians for the dynamic part and $100$k static Gaussians for the static part of the scene, respectively. We perform the same adaptive Gaussian controls for dynamic and static Gaussians as per 3D-GS [40]. Training on a sequence of $300$ frames of $960\times 720$ resolution takes about 2 hours to finish on an A100 GPU. Our rendering FPS is around 40 fps.
4 Experiments
We evaluate our performance quantitatively and qualitatively on a broad range of tasks: long range 3D point tracking, long-range 2D point tracking, and novel view synthesis. We focus our evaluation in particular on datasets that exhibit substantial scene motion. In particular, the iPhone dataset [21] features casual captures of real-world scenes that closely match our target scenarios. It provides comprehensive annotations, including simultaneous validation views, lidar depth, sparse 2D point correspondences across the entire video, and can be used to evaluate our performance on all three tasks. Given the challenge of obtaining precise 3D track annotations for real data, we also evaluate performance using the scenes from the synthetic MOVi-F Kubric dataset [23].
Table 1: Evaluation on iPhone dataset. Our method achieves SOTA performance all tasks of 3D point tracking, 2D point tracking, and novel view synthesis. The baselines that perform best on 2D and 3D tracking (TAPIR [14]+DA [106] and CoTracker [38]+DA [106]) are unable to synthesize novel viewpoints of the scene, while the methods that perform best in novel view synthesis (T-NeRF [21] and HyperNeRF [65]) struggle with or fail to produce 2D and 3D tracks. Our method achieves a significant boost in all three tasks above baselines.
@@ 4 Experiments
[…]
@@ 4.1 Task Specification
Long-range 3D tracking. Our primary task is estimating 3D scene motion for any pixel over long period of time (up to over 10 seconds). For this purpose, we extend the metrics for scene flow evaluation introduced in RAFT-3D [88] to evaluate long-range 3D tracking. Specifically, we report the 3D end-point-error (EPE), which measures the Euclidean distance between the ground truth 3D tracks and predicted tracks at each target time step. In addition, we report the percentage of points that fall within a given threshold of the ground truth 3D location $\delta_{3D}^{.05}=5$cm and $\delta_{3D}^{.10}=10$cm in metric scale.
Long-range 2D tracking. Our globally consistent 3D motion representation can be easily projected onto image plane to get long-range 2D tracks. We therefore also evaluate 2D tracking performance in terms of both position accuracy and occlusion accuracy following the metrics introduced in the TAP-Vid benchmark [13], and report the Average Jaccard (AJ), average position accuracy ($<\delta_{\text{avg}}$), and Occlusion Accuracy (OA).
Novel view synthesis. We measure our method’s novel view synthesis quality as a comprehensive assessment for geometry, appearance, and motion. We evaluate on the iPhone dataset [21] which provides validation views and report co-visibility masked image metrics [21]: mPSNR, mSSIM [96] and mLPIPS [22, 29].
@@
Figure 3: 3D Track visualization on iPhone dataset. We render novel views for each method and overlay their predicted 3D tracks on top of the rendered images. For clarity, we only display a segment of the entire trails spanning 50 frames for a specific set of grid query points. However, it is important to note that our method can generate dense, full-length 3D tracks. Note that TAPIR + DA cannot produce novel view rendering, and we overlay their tracking results with our rendering to make it easier to interpret.
@@
Figure 4: Qualitative comparison of novel-view synthesis on iPhone dataset. The leftmost image in each row shows the training view at the same time step as the validation view. The regions highlighted in green indicate areas excluded from evaluation due to the lack of co-visibility between training and validation views.
@@ 4.2 Baselines
While dynamic novel view synthesis approaches focus primarily on the photometric reconstruction quality and do not explicitly output 3D point tracks, we can adapt some of their representations to derive 3D point tracks for our evaluation. For HyperNeRF [65], we compose the learned inverse mapping (from view space to canonical space) and a forward mapping solved via root-finding [21, 11] to produce 3D tracks at query points. DynIBaR [52] produces short-range view-to-view scene flow, which we chain into long-range 3D tracks for our evaluation. Deformable-3D-GS (D-3DGS) [108] represents dynamic scenes using 3D Gaussians [39] in the canonical space and a deformation MLP network that deforms the canonical 3D Gaussians into each view, which naturally allows 3D motion computation. Finally, T-NeRF [21] models dynamic scenes using time as MLP input in addition to 3D locations, which does not provide a method for extracting 3D motion, and hence they are not considered for 2D/3D tracking evaluation. We evaluate the 3D tracks of these methods only on the iPhone dataset, because we found none can handle the Kubric scene motion.
In addition to dynamic view synthesis baselines, we would also like to include baselines that focus on estimating 3D tracks. However, we did not find prior work that produces long-range 3D tracks from generic monocular videos, hence we adapt existing SOTA methods for long-range 2D tracking and monocular depth estimation. Specifically, we take the state-of-art long-range 2D tracks from TAPIR [14] and CoTracker [38] and lift them into 3D scene motion using monocular depth maps produced by Depth Anything (DA) [106]. We compute the correct scale and shift for each relative depth map from Depth Anything to align them with the scene. The two resulting baselines are called “CoTracker [38] + DA [106]” and “TAPIR [14] + DA [106]”. Note that these baselines can only produce 3D tracks for visible regions, as the depth values for occluded points are unknown from such a 2.5D representation. In contrast, our global representation allows for modeling 3D motion through occlusions.
@@ 4.3 Evaluation on iPhone Dataset
iPhone dataset [21] contains 14 sequences of 200-500 frames featuring various types of challenging real world scene motion. All sequences are recorded using a handheld moving camera in a casual manner, with 7 of them additionally featuring two synchronized static cameras with a large baseline for novel view synthesis evaluation. It also comes with metric depth from the lidar sensors and annotations of 5 to 15 keypoints at ten equally spaced time steps for each sequence. For 3D tracking evaluation, we generate the groundtruth 3D tracks by lifting the 2D keypoint annotations into 3D using lidar depth, and masking out points that are occluded or with invalid lidar depth values.
All experiments are conducted with the original instead of half resolution as in [21] given that our method can handle high-res video input. We also discard Space-out and Wheel scenes due to camera and lidar error. We find that the original camera annotations from ARKit is not accurate enough, and refine them using global bundle adjustment from COLMAP [78].
We report the quantitative comparison in Tab. 4 which shows that our method outperforms all baselines on all tasks by a substantial margin. On 3D and 2D tracking, we show clear improvement over naive baselines of 2D tracks plus depth maps, and significant improvement over all novel view synthesis methods, e.g., nearly halving the EPE of the second best method Deformable-3D-GS [108]. Our method also achieves the best novel view synthesis quality across dynamic NeRF, IBR, and 3D-GS-based baselines.
Fig. 3 shows qualitative comparison of the 3D tracking results. To illustrate 3D tracking quality, we render the novel views and plot the predicted 3D tracks of the given query points onto the novel views. Since “TAPIR + DA” cannot perform novel view synthesis, we overlay their track predictions onto our renderings to aid interpretation. D-3DGS [108] and HyperNeRF [65] fail to capture the significant scene motion in the paper-windmill and spin sequence, resulting in structure degradation and blurry rendering. “TAPIR + DA” can track large motions, but their 3D tracks tend to be noisy and erroneous. In contrast, our method not only generates the highest-quality novel views but also the most smooth and accurate 3D tracks. Fig. 4 provides additional novel view synthesis comparison on the validation views. In Fig. 5, we provide additional visualization of the outputs from our representation. We visualize the rendering of the first three components of PCA decomposition for the motion coefficients, which correlates well with the rigid groups of the moving object.
4.4 Evaluation on the Kubric dataset
Table 2: 3D Tracking evaluation on Kubric dataset.

### v2

@@ 3.3 Optimization
We prepare the following estimates using off-the-shelf methods in our optimization: 1) camera parameters estimated by MegaSaM [62], 2) masks for the moving objects for each frame $\{\mathbf{M}_{t}\}$, which can be easily obtained using Track-Anything [50, 120] with a few user clicks, 3) monocular depth maps $\{\mathbf{D}_{t}\}$ computed using state-of-the-art relative depth estimation method such as Depth Anything [121] and 4) long-range 2D tracks $\{\mathbf{U}_{t\rightarrow t^{\prime}}\}$ for foreground pixels from state-of-the-art point tracking method TAPIR [16]. We align the relative depth maps with the metric depth maps by computing a per-frame global scale and shift and use them for our optimization, as we found relative depth maps tend to contain finer details. We treat the lifted 2D tracks unprojected with the aligned depth maps as noisy initial 3D track observations $\{\mathbf{X}_{t}\}$ for the moving objects. For the static part of the scene, we model them using standard static 3D Gaussians and initialize their 3D locations by unprojecting them into the 3D space using the aligned depth maps. The static and dynamic Gaussians are jointly optimized and rasterized together to form an image. We describe the optimization process below.
@@
Initialization. We first select the canonical frame $t_{0}$ to be the frame in which the most 3D tracks are visible, and initialize the Gaussian means in the canonical frame $\boldsymbol{\mu}_{0}$ as $N$ randomly sampled 3D tracks locations from this set of initial observations. We then perform k-means clustering on the vectorized velocities of the noisy 3D tracks from every frame $\{{\mathbf{X}}_{t}\}$, and initialize the motion bases $\{\mathbf{T}_{0\rightarrow t}^{(b)}\}_{b=1}^{B}$ from these $B$ clusters of tracks. Specifically, for the set of trajectories $\{\mathbf{X}_{t}\}_{b}$ belonging to cluster $b$, we initialize the basis transform ${\mathbf{T}}_{0\rightarrow\tau}^{(b)}$ using weighted Procrustes alignment between the point sets $\{\mathbf{X}_{0}\}_{b}$ and $\{\mathbf{X}_{\tau}\}_{b}$ for all $\tau=0,\dots,T$, where the weights are computed using uncertainty and visibility scores from TAPIR predictions. We initialize $\mathbf{w}^{(b)}$ of each Gaussian to decay exponentially with its distance from the center of cluster $b$ in the canonical frame. We then optimize $\boldsymbol{\mu}_{0}$, $\mathbf{w}^{(b)}$, and set of basis functions $\{\mathbf{T}_{0\rightarrow t}^{(b)}\}_{b=1}^{B}$ to fit the observed 3D tracks with an $\ell_{1}$-loss under temporal smoothness constraints.
@@
Training. We supervise the dynamic Gaussians with two sets of losses. The first set comprise our reconstruction loss to match the per-frame pixelwise color, depth, and masks inputs. The second set enforce consistency of correspondences across time. Specifically, during each training step, we render the image $\hat{\mathbf{I}}_{t}$, depth $\hat{\mathbf{D}}_{t}$, and mask $\hat{\mathbf{M}}_{t}$ from their corresponding training cameras $(\mathbf{K}_{t},\mathbf{E}_{t})$ according to Equation 2. We supervise these predictions with a reconstruction loss applied independently per-frame
$L_{\textrm{recon}}=\|\hat{\mathbf{I}}-\mathbf{I}\|_{1}+\lambda_{\text{depth}}\|\hat{\mathbf{D}}-\mathbf{D}\|_{1}+\lambda_{\textrm{mask}}\|\hat{\mathbf{M}}-\mathbf{M}\|_{1}.$ (7)
@@
The second set of losses supervises the motion of the Gaussians between frames. Specifically, we additionally render the 2D tracks ${\hat{\mathbf{u}}}_{t\rightarrow t^{\prime}}$ and reprojected depths ${\hat{\mathbf{D}}}_{t\rightarrow t^{\prime}}$, for a pair of randomly sampled query time $t$ and target time $t^{\prime}$. We supervise these rendered correspondences with the long-range 2D track estimates via
$L_{\text{track-2d}}=\|\mathbf{U}_{t\rightarrow t^{\prime}}-\hat{\mathbf{U}}_{t\rightarrow t^{\prime}}\|_{1},$ (8)
$L_{\text{track-depth}}=\|\hat{\mathbf{d}}_{t\rightarrow t^{\prime}}-\hat{\mathbf{D}}(\mathbf{U}_{t\rightarrow t^{\prime}})\|_{1}.$ (9)
@@
Finally, we leverage physics motion prior by enforcing a distance-preserving loss between randomly sampled dynamic Gaussians and their k-nearest neighbors. Let $\hat{\mathbf{X}}_{t}$ and $\hat{\mathbf{X}}_{t^{\prime}}$ denote the location of a Gaussian at time $t$ and $t^{\prime}$, and $\mathcal{C}_{k}(\hat{\mathbf{X}}_{t})$ denote the set of k-nearest neighbors of $\hat{\mathbf{X}}_{t}$, we define
@@
Implementation Details. For in-the-wild videos, MegaSaM [62] is used to estimate initial camera poses, which are further refined as learnable parameters during our optimization, updated via the camera gradient function implemented in the gsplat [124] CUDA kernel. For evaluation on public benchmarks, we use the camera annotations provided with each dataset (e.g., from COLMAP [89] or simulation).
We optimize our model using Adam Optimizer [49]. We perform $1000$ iterations of optimization for the initialization phase and $500$ epochs for training phase, respectively. The number of $\mathbb{SE}(3)$ bases $B$ is set to 10 for all of our experiments. We initialize 4040k dynamic Gaussians for the dynamic part and 100100k static Gaussians for the static part of the scene, respectively. We perform the same adaptive Gaussian controls for dynamic and static Gaussians as per 3D-GS [47]. Training on a sequence of $300$ frames of $960\times 720$ resolution takes about 2 hours to finish on an A100 GPU. Our rendering FPS is around 140 fps. We use the same set of hyper-parameters for all experiments.
4 Experiment
We evaluate our performance quantitatively and qualitatively on a broad range of tasks: long range 3D point tracking, long-range 2D point tracking, and novel view synthesis. We focus our evaluation on datasets that exhibit substantial scene motion. In particular, the iPhone dataset [25] features casual captures of real-world scenes that closely match our target scenarios. It provides comprehensive annotations, including simultaneous validation views, lidar depth, sparse 2D point correspondences across the entire video, and can be used to evaluate our performance on all three tasks. Given the challenge of obtaining precise 3D track annotations for real data, we also evaluate performance using the scenes from the synthetic MOVi-F Kubric dataset [28].
@@ 4 Experiment
EPE $\downarrow$ $\delta_{3D}^{.05}\uparrow$ $\delta_{3D}^{.10}\uparrow$ $\text{AJ}\uparrow$ $\hskip-1.99997pt\delta_{\text{avg}}\uparrow$ $\text{OA}\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
@@
Ours + 2DGS[34] 0.097 47.3 71.3 35.8 47.0 87.3 16.75 0.65 0.40
Table 1: Evaluation on iPhone dataset. Our method achieves SOTA performance all tasks of 3D point tracking, 2D point tracking, and novel view synthesis. The baselines that perform best on 2D and 3D tracking (TAPIR [16]+DA [121], CoTracker [45]+DA [121], DELTA [74], SpatialTracker [113]) are unable to synthesize novel views of the scene, while the methods that perform best in novel view synthesis struggle with or fail to produce 2D and 3D tracks. Our method achieves a significant boost in all three tasks above baselines. We include training details about “Ours + 2DGS [34]” in the supplement.
@@ 4.1 Task Specification
Long-Range 3D Tracking. Our primary task is estimating 3D scene motion for any pixel over the entire video. For this purpose, we extend the metrics for scene flow evaluation introduced in RAFT-3D [99] to evaluate long-range 3D tracking. Specifically, we report the 3D end-point-error (EPE), which measures the Euclidean distance between the ground truth 3D tracks and predicted tracks at each target time step. In addition, we report the percentage of points that fall within a given threshold of the ground truth 3D location $\delta_{3D}^{.05}=5$cm and $\delta_{3D}^{.10}=10$cm in metric scale.
Long-Range 2D Tracking. Our 3D motion representation can be easily projected onto image plane to get long-range 2D tracks. Thus we also evaluate 2D tracking performance in terms of both position and occlusion accuracy. Following the metrics introduced in the TAP-Vid benchmark [15], we report the Average Jaccard (AJ), average position accuracy ($\hskip-1.99997pt\delta_{\text{avg}}$), and Occlusion Accuracy (OA).
Novel View Synthesis. We measure our method’s novel view synthesis quality as a comprehensive assessment for geometry, appearance, and motion. We evaluate on the iPhone dataset [25] which provides validation views and report co-visibility masked image metrics [25]: mPSNR, mSSIM [108] and mLPIPS [26, 37]. Additionally, we evaluate on NVIDIA dataset for more comparison.
@@
Figure 3: 3D Track visualization on iPhone dataset. We render novel views for each method and overlay their predicted 3D tracks on top of the rendered images. For clarity, we only display a segment of the trails spanning 50 frames for a specific set of grid query points. However, it is important to note that our method can generate dense, full-length 3D tracks. Note that TAPIR + DA cannot produce novel view rendering, and we overlay their tracking results with our rendering to make it easier to interpret.
@@
Figure 4: Qualitative comparison of novel view synthesis on iPhone dataset. The leftmost image in each row shows the training view at the same time step as the validation view. The regions highlighted in green indicate areas excluded from evaluation due to the lack of co-visibility between training and validation views. Please see the supplemental video for more qualitative results and comparisons.
@@ 4.2 Baselines
While dynamic novel view synthesis approaches focus primarily on the photometric reconstruction quality and do not explicitly output 3D point tracks, we can adapt some of their representations to derive 3D point tracks for our evaluation. For HyperNeRF [77], we compose the learned inverse mapping (from view space to canonical space) and a forward mapping solved via root-finding [25, 11] to produce 3D tracks at query points. DynIBaR [60] produces short-range view-to-view scene flow, which we chain into long-range 3D tracks for our evaluation. Deformable-3D-GS (D-3DGS) [122] represents dynamic scenes using 3D Gaussians [46] in the canonical space and a deformation MLP network that deforms the canonical 3D Gaussians into each view, which naturally allows 3D motion computation. T-NeRF [25] models dynamic scenes using time as MLP input in addition to 3D locations, which does not provide a method for extracting 3D motion, and hence they are not considered for 2D/3D tracking evaluation. Finally, we implemented our own version of DynMF [53] with neural network motion bases. Implementation details are provided in the supplemental material. We evaluate the 3D tracks of these methods only on the iPhone dataset, because we found none can handle the Kubric scene motion.
Alongside dynamic view synthesis baselines, we include 3D tracking baselines. Specifically, we take state-of-the-art long-range 2D tracks from TAPIR [16] and CoTracker [45] and lift them into 3D scene motion using monocular depth maps produced by Depth Anything (DA) [121]. We compute the correct scale and shift for each relative depth map from Depth Anything to align them with the scene. The two resulting baselines are called “CoTracker [45] + DA [121]” and “TAPIR [16] + DA [121]”. Note that these baselines can only produce 3D tracks for visible regions, as the depth values for occluded points are unknown from such a 2.5D representation. In contrast, our global representation allows for modeling 3D motion through occlusions. In addition, we include frame-space 3D tracking methods such as DELTA [74] and SpatialTracker [113], and transform their trajectories into world coordinates using annotated camera poses for evaluation.
@@ 4.3 Evaluation on iPhone Dataset
iPhone dataset [25] contains 14 sequences of 200-500 frames featuring various types of challenging real world scene motion. All sequences are recorded using a handheld moving camera in a casual manner, with 7 of them additionally featuring two synchronized static cameras with a large baseline for novel view synthesis evaluation. For 3D tracking evaluation, we generate the groundtruth 3D tracks by lifting the 2D keypoint annotations into 3D using lidar depth, and masking out points that are occluded or with invalid lidar depth values. All experiments are conducted with the original instead of half resolution as in [25] given that our method can handle high-res video input. We also discard scenes Space-out and Wheel due to camera and lidar error. We find that the original camera annotations from ARKit is not accurate enough, and refine them using global bundle adjustment from COLMAP [88]. The refined poses are used for all methods.
We report the quantitative comparison in Tab. 1 which shows that our method outperforms all baselines on all tasks by a substantial margin. For tracking, we observe significant improvements over naive baselines that combine 2D tracks with depth maps, especially “TAPIR+DA”, which also serves as input to our system, highlighting the effectiveness of our consolidation process. The improvement stems from three factors: 1) a low-dimensional $\mathbb{SE}(3)$ motion representation, which allows inferring occluded motion via nearby visible regions; 2) 3D regularizations, such as low acceleration constraints that promote coherent trajectories; 3) the consolidation of all pairwise tracks into a single 4D representation with persistent 3D geometry, which serves as a structural prior to correct noisy 2D tracks. Our method also achieves the best novel view synthesis quality across dynamic NeRF and 3D-GS-based baselines.
Fig. 3 shows qualitative comparison of the 3D tracking results. We render the novel views and plot the predicted 3D tracks of the given query points onto the novel views. Since “TAPIR + DA” cannot perform novel view synthesis, we overlay their track predictions onto our renderings to aid interpretation. D-3DGS [122] and HyperNeRF [77] fail to capture the significant scene motion in paper-windmill and spin, resulting in structure degradation and blurry rendering. “TAPIR + DA” can track large motions, but their 3D tracks tend to be noisy and erroneous. In contrast, our method not only generates high-quality novel views but also the most smooth and accurate 3D tracks. Fig. 4 provides additional novel view synthesis comparison on the validation views. In Fig. 5, we visualize the rendering of the first three components of PCA decomposition for the motion coefficients, which correlates well with the rigid groups of the moving objects.
4.4 Evaluation on Kubric Dataset